OpenStudy (anonymous):

What polynomial has roots of –1, –2, and 3? a. x^3 + 6x^2 + 11x + 6 b. x^3 – x^2 – 5x + 2 c. x^3 + 2x^2 – 7x + 2 d. x^3 – 7x – 6 Please can u also tell me how?

OpenStudy (auctoratrox):

D. take (x+1)(x+2)(x-3)=...

OpenStudy (callisto):

since it has roots of -1, -2, 3 f(x) = (x+1)(x+2)(x-3) Now, expand the terms, can you do it?

OpenStudy (anonymous):

roots of -1, -2, 3 x=-1 -> (x+1) x=-2 -> (x+2) x=3 -> (x-3) (x+1)(x+2)(x-3) x^3 – 7x – 6 Alternately, you can plug in -1, -2, 3 on all of them and find which one makes all of them 0.

OpenStudy (anonymous):

why do u change -1 to 1 and so on?

jimthompson5910 (jim_thompson5910):

He did this x= - 1, x = -2, or x = 3 x+1 = 0, x+2 = 0, or x -3 = 0 (x+1)(x+2)(x-3) = 0... He used the zero product property here From there, he expanded and simplified the left side

OpenStudy (anonymous):

ok, i get it now... thanks everyone so much!

OpenStudy (anonymous):

i got d... great!